f(x)=h(3x^2+1). Find fَ(1) if hَ(1)=5 and hَ(4)=2.
Answer by ikleyn(52794)

.
f(x) = h(3x^2+1). Find fَ(1) if hَ(1) = 5 and hَ(4) = 2.
~~~~~~~~~~~~~~~~~

They want we find f(1).  So, we substitute x= 1 into f and calculate


    f(1) = by the definition of f(x) = h(3*1^2 + 1) = h(4),



but h(4) is given :  it is 2.   So,


    f(1) = h(4) = 2.    ANSWER.

Solved, answered, explained and completed.

----------------

    The info   " h(1) = 5 "  is an excessive and unnecessary information: it is NOT USED in the solution.

    It is included to the problem by its composer with a unique goal to confuse the reader.